OpenStudy (anonymous):

Assume that the mean length of a human pregnancy is 268 days with a standard deviation of 9 days. What percentage of human pregnancies do we expect to last no more than 260 days?

OpenStudy (anonymous):

\[z=(xbar-\mu)/\sigma\]z=(260-268)/9 z=-0.88888 Use normal distribution tables to find the corresponding percentage: P(X<260) =1- P(Z>-0.8888) P(X<260)=0.187